# Ruiyan Han2∗  Zhen Fang1  Xinyu Sun2∗  Yuchen Ma2  Ziheng Wang2  Yu Zeng122footnotemark: 2 Zehui Chen1   Lin Chen1   Wenxuan Huang3,4   Weijie Xu5   Yi Cao6   Feng Zhao1 1

### v1

@@ D.1 Data Construction
Based on the TIIF benchmark, we generate question–answer pairs for instruction reconstruction, extending the original Text-to-Image (T2I) evaluation to a Text-to-Image-to-Text (T2I2T) setting. To balance task difficulty with answer evaluability, we design task- and question-type–specific prompt templates. For negation-related tasks, we construct prompts that elicit binary (yes/no) questions, ensuring unambiguous evaluation. For tasks like spatial relation, open-ended questions often lead to ambiguous judgments— for instance, an instruction specifies "left" but the generated image places an object in the "front-left" position, an answer such as "in front" may be plausible yet difficult to assess consistently. To improve evaluation stability, we therefore formulate these tasks as multiple-choice questions. For the remaining task types, such as color recognition and counting, we adopt open-ended question–answer formats to maintain sufficient difficulty and discriminative power. Moreover, we explicitly enforce task-type–based question completeness: since all instruction-implied information relevant to the task type is treated as a reconstruction target, a QA set is considered valid only if it fully covers the reconstruction targets without redundancy. QA pairs are generated using Qwen3-235B-A22B yang2025qwen3 and subsequently annotated under a unified labeling protocol by experienced human annotators. After filtering, we retain 1,401 high-quality instances and totally 2968 questions ( The distribution of question types is shown in Tab.16) for evaluation, covering almost all task types present in the original TIIF benchmark. We present several cases in Fig. 10.
@@ E.1 Quantitative Experiments
Detailed scores across the six T2I benchmarks are reported in Tab. 9,  10,  111213and  14. We also evaluate our model on the image edit task rise in Tab. 6.
@@ E.3 Failure Cases
Janus-Pro (januspro2025) 0.553 0.001 0.139 0.276 0.365 0.267
T2I-R1 (jiang2025t2i) 0.804 0.073 0.167 0.290 0.277 0.322
BLIP3-o (chen2025blip3) 0.711 0.013 0.223 0.361 0.229 0.307
BAGEL (deng2025emerging) 0.769 0.244 0.173 0.367 0.251 0.361
Show-o2-7B (xie2025show) 0.817 0.002 0.226 0.317 0.177 0.308
SDv1.5 (rombach2022high) 0.565 0.010 0.207 0.383 0.429 0.319
SDXL (sdxl) 0.688 0.029 0.237 0.332 0.296 0.316
FLUX.1-dev (flux) 0.786 0.523 0.253 0.368 0.238 0.434
SD3 (SD3) 0.805 0.407 0.293 0.386 0.244 0.427
FLUX.1-dev (flux) 0.786 0.523 0.253 0.368 0.238 0.434
SANA-1.5 4.8B (PAG) (xie2025sana) 0.765 0.069 0.217 0.401 0.216 0.334
Lumina-Image 2.0 (qin2025lumina) 0.819 0.106 0.270 0.354 0.216 0.353
IRG*  (huang2025interleaving) 0.839 0.377 0.239 0.427 0.192 0.415
OmniGen2  (xiao2025omnigen) 0.804 0.680 0.271 0.377 0.242 0.475
@@
FLUX.1-dev (flux) 66.24 66.72 74.41 76.67 72.50 75.50 78.20 79.78 72.52 74.73 60.72 60.95 66.76 65.50 61.76 60.74 56.60 57.49 63.33 60.00 44.49 54.75 74.63 72.01
FLUX.1-Pro (flux) 63.75 63.53 71.39 73.57 70.00 68.50 68.51 79.97 75.66 72.23 64.63 61.42 70.69 72.99 62.34 57.27 64.65 57.11 63.00 63.00 34.39 36.65 69.94 66.78
DALL-E 3 (dalle3) 74.47 72.94 77.35 78.40 77.62 75.00 80.22 79.67 74.22 80.54 70.11 68.45 76.65 75.05 68.39 68.07 63.64 59.92 79.31 80.00 74.07 75.51 76.12 62.69
SD3.5-large (SD3) 68.69 64.92 73.72 72.10 77.50 66.50 74.79 77.16 68.87 72.64 65.59 63.41 70.85 68.22 65.03 62.93 61.03 61.66 56.67 60.00 73.30 46.15 70.15 69.03
PixArt-$\Sigma$ (chen2024pixart) 57.46 57.04 67.74 68.19 65.50 69.50 74.33 72.11 63.40 62.96 56.71 54.52 62.47 59.67 57.51 55.08 54.84 52.64 76.67 73.33 2.71 4.98 63.06 63.06
Show-o (xie2024show) 57.34 61.33 69.99 75.30 66.50 80.00 76.47 71.88 67.00 74.04 58.25 58.19 67.21 64.33 54.26 58.86 61.38 56.19 46.67 66.67 4.98 11.31 71.64 68.66
Janus-Pro-7B (januspro2025) 65.38 61.10 74.99 73.19 74.50 78.00 73.69 70.51 76.77 71.04 61.77 56.03 65.71 66.48 62.01 55.62 61.16 49.34 43.33 70.00 38.46 42.08 79.48 73.51
T2I-R1 (jiang2025t2i) 67.61 68.34 81.14 79.45 80.50 78.50 83.09 79.49 79.81 80.37 67.38 65.90 69.92 65.27 70.10 71.62 68.69 64.68 50.00 63.33 32.13 37.56 74.25 74.25
BAGEL (deng2025emerging) 70.97 71.79 78.16 78.12 78.00 79.50 80.24 79.08 76.25 75.77 68.23 68.19 73.37 77.49 64.36 66.15 68.92 61.48 80.00 80.00 40.72 52.40 76.87 74.63
MidJourney v7 (midjourneyv7) 65.92 62.43 73.96 74.63 75.00 82.00 78.74 78.51 68.12 68.55 63.44 62.59 70.60 74.03 64.43 59.58 58.84 61.34 66.67 33.33 31.67 34.39 79.22 75.32
Show-o2  (xie2025show) 62.80 63.87 75.30 74.45 73.00 71.00 77.22 74.09 75.69 78.25 61.38 66.12 63.47 67.44 62.63 70.31 64.15 60.00 60.00 33.33 14.03 10.86 75.00 74.63
BAGEL (deng2025emerging) 68.06 68.78 77.63 79.40 75.00 77.00 78.55 82.37 79.33 78.81 71.24 68.20 77.65 75.37 69.77 65.87 72.93 67.91 69.93 63.33 26.24 26.70 69.78 71.64
IRG*  (huang2025interleaving) 76.00 73.77 83.17 81.28 81.00 76.00 82.96 81.86 85.54 85.98 75.25 74.66 75.82 77.25 78.16 77.76 73.84 72.93 90.00 70.00 43.89 47.51 72.76 74.63
@@
Gen. Only SDv1.5 (rombach2022high) 0.34 0.35 0.32 0.28 0.29 0.21 0.32
SDXL (sdxl) 0.43 0.48 0.47 0.44 0.45 0.27 0.43
SD3.5-large (SD3) 0.44 0.50 0.58 0.44 0.52 0.31 0.46
PixArt-Alpha (chen2024pixart) 0.45 0.50 0.48 0.49 0.56 0.34 0.47
playground-v2.5 (li2024playground) 0.49 0.58 0.55 0.43 0.48 0.33 0.49
FLUX.1-dev (flux) 0.48 0.58 0.62 0.42 0.51 0.35 0.50
Unified Janus (wu2025janus) 0.16 0.26 0.35 0.28 0.30 0.14 0.23
Show-o-512 (xie2024show) 0.28 0.40 0.48 0.30 0.46 0.30 0.35
Janus-Pro-7B (januspro2025) 0.30 0.37 0.49 0.36 0.42 0.26 0.35
Emu3 (emu3) 0.34 0.45 0.48 0.41 0.45 0.27 0.39
MetaQuery-XL (pan2025transfer) 0.56 0.55 0.62 0.49 0.63 0.41 0.55
BAGEL (deng2025emerging) 0.42 0.53 0.64 0.42 0.57 0.43 0.50
Show-o2 (xie2025show) 0.64 0.58 0.61 0.58 0.63 0.49 0.61
T2I-R1  (jiang2025t2i) 0.56 0.55 0.63 0.54 0.55 0.30 0.54
BLIP3-o  (blip3o) 0.49 0.51 0.63 0.54 0.63 0.37 0.52
@@
GPT-4o (openai2025chatgpt4o) 0.81 0.71 0.89 0.83 0.79 0.74 0.80
Table 12: Comprehensive T2I-CompBench Results. This table includes T2I (flux; SD3; sdxl) and UMMs (januspro2025; xie2025show).
@@
Gen. Only PixArt-$\alpha$ (chen2024pixart) 0.98 0.50 0.44 0.80 0.08 0.07 0.48
SDv2.12.1 (rombach2022high) 0.98 0.51 0.44 0.85 0.07 0.17 0.50
DALL-E $2$ (dalle2) 0.94 0.66 0.49 0.77 0.10 0.19 0.52
Emu33-Gen  (emu3) 0.98 0.71 0.34 0.81 0.17 0.21 0.54
SDXL (sdxl) 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E $3$ (dalle3) 0.96 0.87 0.47 0.83 0.43 0.45 0.67
SD3-Medium (SD3) 0.99 0.94 0.72 0.89 0.33 0.60 0.74
FLUX.1-dev (flux) 0.98 0.93 0.75 0.93 0.68 0.65 0.82
Unified Chameleon (chameleon) - - - - - - 0.39
LWM (lwm) 0.93 0.41 0.46 0.79 0.09 0.15 0.47
SEED-X (seed-x) 0.97 0.58 0.26 0.80 0.19 0.14 0.49
TokenFlow-XL (qu2024tokenflow) 0.95 0.60 0.41 0.81 0.16 0.24 0.55
ILLUME (wang2024illume) 0.99 0.86 0.45 0.71 0.39 0.28 0.61
Janus (wu2025janus) 0.97 0.68 0.30 0.84 0.46 0.42 0.61
Transfusion (transfusion) - - - - - - 0.63
Emu33-Gen(emu3) 0.99 0.81 0.42 0.80 0.49 0.45 0.66
Show-o (xie2024show) 0.98 0.80 0.66 0.84 0.31 0.50 0.68
Janus-Pro-7B (januspro2025) 0.99 0.89 0.59 0.90 0.79 0.66 0.80
MetaQuery-XL (pan2025transfer) - - - - - - 0.80
BAGEL (deng2025emerging) 0.99 0.95 0.76 0.87 0.50 0.60 0.78
Show-o2 (xie2025show) 1.00 0.87 0.58 0.92 0.52 0.62 0.76
BAGEL (deng2025emerging) 0.99 0.92 0.75 0.89 0.54 0.63 0.79
IRG*  (huang2025interleaving) 0.98 0.94 0.83 0.86 0.74 0.73 0.85
UniGen*  (tian2025unigen) 1.00 0.94 0.78 0.87 0.57 0.54 0.78
UniRL  (unirl) 0.96 0.80 0.67 0.86 0.50 0.67 0.74
@@
PixArt-$\alpha$ (chen2024pixart) 74.97 79.32 78.60 82.57 76.96 71.11
Lumina-Next (zhuo2024luminanext) 82.82 88.65 86.44 80.53 81.82 74.63
Playground v2.5 (li2024playground) 83.06 82.59 81.20 84.08 83.50 75.47
Hunyuan-DiT (li2024hunyuandit) 84.59 80.59 88.01 74.36 86.41 78.87
Janus (wu2025janus) 82.33 87.38 87.70 85.46 86.41 79.68
Janus-Pro-1B (chen2025janus) 87.58 88.63 88.17 88.98 88.30 82.63
DALL-E 3 (dalle3) 90.97 89.61 88.39 90.58 89.83 83.50
FLUX.1-dev  (flux) 74.35 90.00 88.96 90.87 88.33 83.84
SD3 Medium (SD3) 87.90 91.01 88.83 80.70 88.68 84.08
Janus-Pro-7B (chen2025janus) 86.90 88.90 89.40 89.32 89.48 84.19
BAGEL  (deng2025emerging) - - - - - 84.03
UniCorn  (xie2025reconstruction) 91.62 91.97 91.39 91.22 91.64 86.83

### v2

@@ D.1 Data Construction
Based on the TIIF benchmark, we generate question–answer pairs for instruction reconstruction, extending the original Text-to-Image (T2I) evaluation to a Text-to-Image-to-Text (T2I2T) setting. To balance task difficulty with answer evaluability, we design task- and question-type–specific prompt templates. For negation-related tasks, we construct prompts that elicit binary (yes/no) questions, ensuring unambiguous evaluation. For tasks like spatial relation, open-ended questions often lead to ambiguous judgments— for instance, an instruction specifies "left" but the generated image places an object in the "front-left" position, an answer such as "in front" may be plausible yet difficult to assess consistently. To improve evaluation stability, we therefore formulate these tasks as multiple-choice questions. For the remaining task types, such as color recognition and counting, we adopt open-ended question–answer formats to maintain sufficient difficulty and discriminative power. Moreover, we explicitly enforce task-type–based question completeness: since all instruction-implied information relevant to the task type is treated as a reconstruction target, a QA set is considered valid only if it fully covers the reconstruction targets without redundancy. QA pairs are generated using Qwen3-235B-A22B Yang et al. (2025) and subsequently annotated under a unified labeling protocol by experienced human annotators. After filtering, we retain 1,401 high-quality instances and totally 2968 questions ( The distribution of question types is shown in Tab.16) for evaluation, covering almost all task types present in the original TIIF benchmark. We present several cases in Fig. 10.
@@ E.1 Quantitative Experiments
Detailed scores across the six T2I benchmarks are reported in Tab. 9,  10,  111213and  14. We also evaluate our model on the image edit task Zhao et al. (2025b) in Tab. 6.
@@ E.3 Failure Cases
Janus-Pro (Chen et al., 2025d) 0.553 0.001 0.139 0.276 0.365 0.267
T2I-R1 (Jiang et al., 2025) 0.804 0.073 0.167 0.290 0.277 0.322
BLIP3-o (Chen et al., 2025c) 0.711 0.013 0.223 0.361 0.229 0.307
BAGEL (Deng et al., 2025a) 0.769 0.244 0.173 0.367 0.251 0.361
Show-o2-7B (Xie et al., 2025c) 0.817 0.002 0.226 0.317 0.177 0.308
SDv1.5 (Rombach et al., 2022) 0.565 0.010 0.207 0.383 0.429 0.319
SDXL (Podell et al., 2024) 0.688 0.029 0.237 0.332 0.296 0.316
FLUX.1-dev (Labs, 2024) 0.786 0.523 0.253 0.368 0.238 0.434
SD3 (Esser et al., 2024) 0.805 0.407 0.293 0.386 0.244 0.427
FLUX.1-dev (Labs, 2024) 0.786 0.523 0.253 0.368 0.238 0.434
SANA-1.5 4.8B (PAG) (Xie et al., 2025a) 0.765 0.069 0.217 0.401 0.216 0.334
Lumina-Image 2.0 (Qin et al., 2025) 0.819 0.106 0.270 0.354 0.216 0.353
IRG*  (Huang et al., 2025c) 0.839 0.377 0.239 0.427 0.192 0.415
OmniGen2  (Xiao et al., 2025) 0.804 0.680 0.271 0.377 0.242 0.475
@@
FLUX.1-dev (Labs, 2024) 66.24 66.72 74.41 76.67 72.50 75.50 78.20 79.78 72.52 74.73 60.72 60.95 66.76 65.50 61.76 60.74 56.60 57.49 63.33 60.00 44.49 54.75 74.63 72.01
FLUX.1-Pro (Labs, 2024) 63.75 63.53 71.39 73.57 70.00 68.50 68.51 79.97 75.66 72.23 64.63 61.42 70.69 72.99 62.34 57.27 64.65 57.11 63.00 63.00 34.39 36.65 69.94 66.78
DALL-E 3 (Betker et al., 2023b) 74.47 72.94 77.35 78.40 77.62 75.00 80.22 79.67 74.22 80.54 70.11 68.45 76.65 75.05 68.39 68.07 63.64 59.92 79.31 80.00 74.07 75.51 76.12 62.69
SD3.5-large (Esser et al., 2024) 68.69 64.92 73.72 72.10 77.50 66.50 74.79 77.16 68.87 72.64 65.59 63.41 70.85 68.22 65.03 62.93 61.03 61.66 56.67 60.00 73.30 46.15 70.15 69.03
PixArt-$\Sigma$ (Chen et al., 2024b) 57.46 57.04 67.74 68.19 65.50 69.50 74.33 72.11 63.40 62.96 56.71 54.52 62.47 59.67 57.51 55.08 54.84 52.64 76.67 73.33 2.71 4.98 63.06 63.06
Show-o (Xie et al., 2024) 57.34 61.33 69.99 75.30 66.50 80.00 76.47 71.88 67.00 74.04 58.25 58.19 67.21 64.33 54.26 58.86 61.38 56.19 46.67 66.67 4.98 11.31 71.64 68.66
Janus-Pro-7B (Chen et al., 2025d) 65.38 61.10 74.99 73.19 74.50 78.00 73.69 70.51 76.77 71.04 61.77 56.03 65.71 66.48 62.01 55.62 61.16 49.34 43.33 70.00 38.46 42.08 79.48 73.51
T2I-R1 (Jiang et al., 2025) 67.61 68.34 81.14 79.45 80.50 78.50 83.09 79.49 79.81 80.37 67.38 65.90 69.92 65.27 70.10 71.62 68.69 64.68 50.00 63.33 32.13 37.56 74.25 74.25
BAGEL (Deng et al., 2025a) 70.97 71.79 78.16 78.12 78.00 79.50 80.24 79.08 76.25 75.77 68.23 68.19 73.37 77.49 64.36 66.15 68.92 61.48 80.00 80.00 40.72 52.40 76.87 74.63
MidJourney v7 (Midjourney, 2025) 65.92 62.43 73.96 74.63 75.00 82.00 78.74 78.51 68.12 68.55 63.44 62.59 70.60 74.03 64.43 59.58 58.84 61.34 66.67 33.33 31.67 34.39 79.22 75.32
Show-o2  (Xie et al., 2025c) 62.80 63.87 75.30 74.45 73.00 71.00 77.22 74.09 75.69 78.25 61.38 66.12 63.47 67.44 62.63 70.31 64.15 60.00 60.00 33.33 14.03 10.86 75.00 74.63
BAGEL (Deng et al., 2025a) 68.06 68.78 77.63 79.40 75.00 77.00 78.55 82.37 79.33 78.81 71.24 68.20 77.65 75.37 69.77 65.87 72.93 67.91 69.93 63.33 26.24 26.70 69.78 71.64
IRG*  (Huang et al., 2025c) 76.00 73.77 83.17 81.28 81.00 76.00 82.96 81.86 85.54 85.98 75.25 74.66 75.82 77.25 78.16 77.76 73.84 72.93 90.00 70.00 43.89 47.51 72.76 74.63
@@
Gen. Only SDv1.5 (Rombach et al., 2022) 0.34 0.35 0.32 0.28 0.29 0.21 0.32
SDXL (Podell et al., 2024) 0.43 0.48 0.47 0.44 0.45 0.27 0.43
SD3.5-large (Esser et al., 2024) 0.44 0.50 0.58 0.44 0.52 0.31 0.46
PixArt-Alpha (Chen et al., 2024b) 0.45 0.50 0.48 0.49 0.56 0.34 0.47
playground-v2.5 (Li et al., 2024a) 0.49 0.58 0.55 0.43 0.48 0.33 0.49
FLUX.1-dev (Labs, 2024) 0.48 0.58 0.62 0.42 0.51 0.35 0.50
Unified Janus (Wu et al., 2025b) 0.16 0.26 0.35 0.28 0.30 0.14 0.23
Show-o-512 (Xie et al., 2024) 0.28 0.40 0.48 0.30 0.46 0.30 0.35
Janus-Pro-7B (Chen et al., 2025d) 0.30 0.37 0.49 0.36 0.42 0.26 0.35
Emu3 (Wang et al., 2024b) 0.34 0.45 0.48 0.41 0.45 0.27 0.39
MetaQuery-XL (Pan et al., 2025) 0.56 0.55 0.62 0.49 0.63 0.41 0.55
BAGEL (Deng et al., 2025a) 0.42 0.53 0.64 0.42 0.57 0.43 0.50
Show-o2 (Xie et al., 2025c) 0.64 0.58 0.61 0.58 0.63 0.49 0.61
T2I-R1  (Jiang et al., 2025) 0.56 0.55 0.63 0.54 0.55 0.30 0.54
BLIP3-o  (Chen et al., 2025b) 0.49 0.51 0.63 0.54 0.63 0.37 0.52
@@
GPT-4o (OpenAI, 2025) 0.81 0.71 0.89 0.83 0.79 0.74 0.80
Table 12: Comprehensive T2I-CompBench Results. This table includes T2I (Labs, 2024; Esser et al., 2024; Podell et al., 2024) and UMMs (Chen et al., 2025d; Xie et al., 2025c).
@@
Gen. Only PixArt-$\alpha$ (Chen et al., 2024b) 0.98 0.50 0.44 0.80 0.08 0.07 0.48
SDv2.12.1 (Rombach et al., 2022) 0.98 0.51 0.44 0.85 0.07 0.17 0.50
DALL-E $2$ (Ramesh et al., 2022) 0.94 0.66 0.49 0.77 0.10 0.19 0.52
Emu33-Gen  (Wang et al., 2024b) 0.98 0.71 0.34 0.81 0.17 0.21 0.54
SDXL (Podell et al., 2024) 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E $3$ (Betker et al., 2023b) 0.96 0.87 0.47 0.83 0.43 0.45 0.67
SD3-Medium (Esser et al., 2024) 0.99 0.94 0.72 0.89 0.33 0.60 0.74
FLUX.1-dev (Labs, 2024) 0.98 0.93 0.75 0.93 0.68 0.65 0.82
Unified Chameleon (Team, 2024) - - - - - - 0.39
LWM (Liu et al., 2024a) 0.93 0.41 0.46 0.79 0.09 0.15 0.47
SEED-X (Ge et al., 2024) 0.97 0.58 0.26 0.80 0.19 0.14 0.49
TokenFlow-XL (Qu et al., 2024) 0.95 0.60 0.41 0.81 0.16 0.24 0.55
ILLUME (Wang et al., 2024a) 0.99 0.86 0.45 0.71 0.39 0.28 0.61
Janus (Wu et al., 2025b) 0.97 0.68 0.30 0.84 0.46 0.42 0.61
Transfusion (Zhou et al., 2024a) - - - - - - 0.63
Emu33-Gen(Wang et al., 2024b) 0.99 0.81 0.42 0.80 0.49 0.45 0.66
Show-o (Xie et al., 2024) 0.98 0.80 0.66 0.84 0.31 0.50 0.68
Janus-Pro-7B (Chen et al., 2025d) 0.99 0.89 0.59 0.90 0.79 0.66 0.80
MetaQuery-XL (Pan et al., 2025) - - - - - - 0.80
BAGEL (Deng et al., 2025a) 0.99 0.95 0.76 0.87 0.50 0.60 0.78
Show-o2 (Xie et al., 2025c) 1.00 0.87 0.58 0.92 0.52 0.62 0.76
BAGEL (Deng et al., 2025a) 0.99 0.92 0.75 0.89 0.54 0.63 0.79
IRG*  (Huang et al., 2025c) 0.98 0.94 0.83 0.86 0.74 0.73 0.85
UniGen*  (Tian et al., 2025) 1.00 0.94 0.78 0.87 0.57 0.54 0.78
UniRL  (Mao et al., 2025) 0.96 0.80 0.67 0.86 0.50 0.67 0.74
@@
PixArt-$\alpha$ (Chen et al., 2024b) 74.97 79.32 78.60 82.57 76.96 71.11
Lumina-Next (Zhuo et al., 2024) 82.82 88.65 86.44 80.53 81.82 74.63
Playground v2.5 (Li et al., 2024a) 83.06 82.59 81.20 84.08 83.50 75.47
Hunyuan-DiT (Li et al., 2024c) 84.59 80.59 88.01 74.36 86.41 78.87
Janus (Wu et al., 2025b) 82.33 87.38 87.70 85.46 86.41 79.68
Janus-Pro-1B (Chen et al., 2025e) 87.58 88.63 88.17 88.98 88.30 82.63
DALL-E 3 (Betker et al., 2023b) 90.97 89.61 88.39 90.58 89.83 83.50
FLUX.1-dev  (Labs, 2024) 74.35 90.00 88.96 90.87 88.33 83.84
SD3 Medium (Esser et al., 2024) 87.90 91.01 88.83 80.70 88.68 84.08
Janus-Pro-7B (Chen et al., 2025e) 86.90 88.90 89.40 89.32 89.48 84.19
BAGEL  (Deng et al., 2025a) - - - - - 84.03
UniCorn 91.62 91.97 91.39 91.22 91.64 86.83